# Learning to Animate Images from A Few Videos to Portray Delicate Human Actions

### v1

@@ 4.2 Ablation Studies
Joint Training with Multiple Action Classes. We examine whether the model benefits from joint training across multiple action classes. We use the training videos from the four action classes (sprint run, soccer shoot, canoeing sprint, and hugging human) to train a single model. The results in Table A1 show improvements across nearly all metrics. The improvements in Image Alignment, Temporal Consistency, and Cosine RGB are considerable. The results suggest that joint training with multiple action classes enhances the quality of the generated videos. This makes our technique more practical for applications that need to animate images to portray multiple delicate or customized human actions.
@@
Analysis of Motion Alignment Module. In Table A2, we compare the performance of different $\tau$ values in Variant #3 and different $p$ values in Variant #4. For $\tau$, we observe that decreasing $\tau$ reduces performance in Temporal Consistency, CD-FVD, and Cosine Flow, especially in Temporal Consistency (94.85 for $\tau=25$) and Cosine Flow (0.6979 for $\tau=25$). This suggests that including more channels in motion features degrades video quality, likely because motion information is only encoded in a limited number of channels [71]. Thus, we set $\tau=90$ for the remaining experiments. Regarding $p$, substituting inter-frame correspondence relations in all cross-frame attention layers ($p=1.0$) lowers Cosine RGB and Cosine Flow (e.g., Cosine Flow drops to 0.6934 for $p=1.0$). This might be due to the excessive regularization from substituting inter-frame correspondence relations in every layer, which makes learning difficult. Therefore, we use $p=0.5$ in the remaining experiments.
@@
Analysis of Detail Enhancement Decoder. In Table A2, we compare the effects of the Warping Branch and the Patch Attention Branch in Variant #6. Using only the Warping Branch leads to a notable improvement in Temporal Consistency (from 95.01 to 95.35). In contrast, the Patch Attention Branch provides a modest increase in Text Alignment (from 22.52 to 22.71) but results in a significant drop in Image Alignment (from 76.35 to 74.97). When both branches are combined, there is an enhancement in both Text Alignment and Temporal Consistency, accompanied by only a slight decrease in Image Alignment. These results suggest that the two branches have complementary effects.
@@
To evaluate the effectiveness of FLASH on additional datasets, we conducted experiments on the UCF Sports Action Dataset [59], focusing on two actions: golf swing and lifting. Due to the limited number of videos in this dataset, only 6 videos (golf swing) and 4 videos (lifting) are used for training.
@@
Table A3: Quantitative comparison of different methods on the UCF Sports Action Dataset.
Method Text Alignment ($\uparrow$) Image Alignment ($\uparrow$) Temporal Consistency ($\uparrow$)   CD-FVD ($\downarrow$)    Cosine RGB ($\uparrow$)    Cosine Flow ($\uparrow$)
TI2V-Zero 24.62 70.16 88.87 1222.35 0.7190 0.6443
SparseCtrl 23.26 61.16 89.69 1574.69 0.7156 0.6321
PIA 23.93 66.03 94.41 1385.54 0.7005 0.5851
DynamiCrafter 23.95 87.84 96.75 1630.83 0.7783 0.6316
DreamVideo 26.04 78.20 96.19 949.72 0.6860 0.7020
LAMP 24.02 81.17 95.17 1076.77 0.7415 0.7378
FLASH 24.11 85.75 96.22 480.70 0.8680 0.7936
@@
Although FLASH can animate diverse reference images, it encounters challenges in accurately generating interactions involving human objects, particularly when multiple objects are present. For example, in Figure A6 (a), while a chopping action is depicted, the object being chopped is not the wood. Furthermore, if the initial action states in the reference images differ noticeably in motion details from those in the training videos, the model may struggle with animation. For example, in Figure A6 (b), the initial action status suggests a small-scale motion for chopping wood, which differs from the large-scale motion in training videos; in Figure A6 (c), the knee elevation motion contrasts with the steadier motion of running in place observed in the training videos; and in Figure A6 (d), a baby holding a cup with both hands deviates from the adult actions in the training videos, where one hand is used to hold the cup while drinking water. These results suggest that the model still lacks a thorough understanding of motion and interactions. Leveraging advanced multi-modal large language models to improve the understanding of human-object interactions could be a promising approach to addressing these challenges.
@@
We firmly oppose the misuse of generative AI for creating harmful content or spreading false information. We do not assume any responsibility for potential misuse by users. Nonetheless, we recognize that our approach, which focuses on animation human images, carries the risk of potential misuse. To address these risks, we are committed to maintaining the highest ethical standards in our research by complying with legal requirements and protecting privacy. Additionally, we will explore implementing a content safety mechanism, similar to the one used in Stable Diffusion [54], as an effective way to address these concerns.
A7 Conclusion
In this paper, we introduce FLASH, a model that learns to animate images to depict human actions using minimal training data. We propose the Motion Alignment Module to learn consistent motion signals between videos with identical motion but different appearances, facilitating the learning of generalizable motion patterns. Additionally, we devise the Detail Enhancement Decoder to enhance details in generated frames, improving transition smoothness. Experimental results show that FLASH effectively animates diverse images with unseen actor or scene appearances, including those with unrealistic scenarios and humanoid aliens, into specified actions while maintaining smooth transitions.

### v2

@@ 5 Conclusion
Joint Training with Multiple Action Classes. We examine whether the model benefits from joint training across multiple action classes. We use all the training videos from the four action classes (sprint run, soccer shoot, canoeing sprint, and hugging human) to train a single model. The results in Table A1 show improvements across all metrics. The improvements in Image Alignment, Temporal Consistency, and Cosine RGB are considerable. The results suggest that joint training with multiple action classes enhances the quality of the generated videos. This makes our technique more practical for applications that have accessible example videos of multiple delicate or customized human actions.
@@
Analysis of Motion Alignment Module. In Table A2, we compare the performance of different $\tau$ values in Variant #3 and different $p$ values in Variant #4. For $\tau$, we observe that decreasing $\tau$ reduces performance in Temporal Consistency, CD-FVD, and Cosine Flow, especially in Temporal Consistency (94.85 for $\tau=25$) and Cosine Flow (69.79 for $\tau=25$). This suggests that including more channels as motion channels degrades video quality, likely because motion information is only encoded in a limited number of channels [93], and aligning too many channels hampers feature learning. Thus, we set $\tau=90$ for the remaining experiments. Regarding $p$, substituting inter-frame correspondence relations in all cross-frame attention layers ($p=1.0$) lowers Cosine Flow significantly (e.g., 69.34 for $p=1.0$) but doesn’t affect other metrics obviously. This might be due to the excessive regularization from substituting inter-frame correspondence relations in every layer, which makes learning difficult. Therefore, we use $p=0.5$ in the remaining experiments.
@@
Analysis of Detail Enhancement Decoder. In Table A2, we compare the effects of the Warping Branch and the Patch Attention Branch in Variant #6. Using only the Warping Branch leads to a notable improvement in Temporal Consistency (from 95.01 to 95.35). In contrast, the Patch Attention Branch provides a modest increase in Text Alignment (from 22.52 to 22.71) but results in a significant drop in Image Alignment (from 76.35 to 74.97). When both branches are combined, there is an enhancement in both Text Alignment and Temporal Consistency, accompanied by only a slight decrease in Image Alignment. These results suggest that the two branches have complementary effects. Therefore, we use the two branches in the Detail Enhancement Decoder.
Evaluation of the Detail Enhancement Decoder with DINO. The CLIP vision encoder, trained on vision-language tasks, may have limited ability to perceive fine-grained visual details [81], which can affect the evaluation of Image Alignment and Temporal Consistency. Therefore, we use the DINO-V2 [62] vision encoder, which excels at capturing rich, fine-grained details at the pixel level, to assess Image Alignment and Temporal Consistency. The results in Table A3 demonstrate that the Detail Enhancement Decoder enhances both Image Alignment and Temporal Consistency, illustrating its effectiveness in improving transition smoothness.
Table A4: Quantitative comparison of different methods on the UCF Sports Action Dataset. The best and second-best results are bolded and underlined.
Method    Cosine RGB ($\uparrow$)    Cosine Flow ($\uparrow$)   CD-FVD ($\downarrow$) Text Alignment ($\uparrow$) Image Alignment ($\uparrow$) Temporal Consistency ($\uparrow$)
TI2V-Zero 71.90 64.43 1222.35 24.62 70.16 88.87
SparseCtrl 71.56 63.21 1574.69 23.26 61.16 89.69
PIA 70.05 58.51 1385.54 23.93 66.03 94.41
DynamiCrafter 77.83 63.16 1630.83 23.95 87.84 96.75
DreamVideo 68.60 70.20 949.72 26.04 78.20 96.19
MotionDirector 75.60 63.01 1315.36 23.88 76.92 97.07
LAMP 74.15 73.78 1076.77 24.02 81.17 95.17
FLASH 86.80 79.36 480.70 24.11 85.75 96.22
Table A5: Quantitative comparison of different methods on non-human motion videos. The best and second-best results are bolded and underlined.
Method    Cosine RGB ($\uparrow$)    Cosine Flow ($\uparrow$)   CD-FVD ($\downarrow$) Text Alignment ($\uparrow$) Image Alignment ($\uparrow$) Temporal Consistency ($\uparrow$)
TI2V-Zero 58.96 45.31 1562.76 21.95 79.05 93.00
SparseCtrl 67.89 59.02 1441.31 21.92 76.89 93.75
PIA 68.11 57.02 1591.11 21.83 79.44 96.83
DynamiCrafter 77.14 69.90 1371.39 22.18 87.27 98.08
DreamVideo 68.80 61.44 1222.22 22.75 84.40 96.96
MotionDirector 74.57 68.41 1302.02 20.75 79.09 96.68
LAMP 79.48 71.89 1210.55 22.14 86.68 97.49
FLASH 79.53 75.48 1204.72 22.05 85.42 97.51
Figure A6: Failure cases of FLASH.
@@
To evaluate the effectiveness of FLASH on additional datasets, we conducted experiments on the UCF Sports Action Dataset [79], focusing on two actions: golf swing and lifting. Due to the limited number of videos in this dataset, we use only 6 golf swing videos and 4 lifting videos for training. For each class, we use the first frames of two videos for testing.
@@
A4.4 Experiments on Non-human Motion Videos
To assess the performance of FLASH on non-human motion videos, we conducted experiments using two categories of natural motion: firework and raining. The videos were sourced from [92]. For each category, we selected two videos for testing and used the remaining videos for training.
Table A5 presents a comparison between FLASH and baseline methods. FLASH achieves superior performance in CD-FVD, Cosine RGB, and Cosine Flow while not showing a obvious decline in CLIP scores. Although DynamiCrafter performs better in Image Alignment and Temporal Consistency, it struggles with CD-FVD and Cosine Flow. Similarly, DreamVideo excels in Text Alignment but performs poorly in Cosine RGB and Cosine Flow. These results indicate that FLASH can also animate images into videos depicting natural scene motion.
@@
Although FLASH can animate diverse reference images, it encounters challenges in accurately generating interactions involving human and objects, particularly when multiple objects are present. For example, in Figure A6(a), while a chopping action is depicted, the object being chopped is not the wood. Furthermore, if the initial action states in the reference images differ noticeably in motion patterns from those in the training videos, the model may struggle with animation. For example, in Figure A6(b), the initial action status suggests a small-scale motion for chopping wood, which differs from the large-scale motion in training videos; in Figure A6(c), the knee elevation motion contrasts with the steadier motion of running in place observed in the training videos; and in Figure A6(d), a baby holding a cup with both hands deviates from the adult actions in the training videos, where one hand is used to hold the cup while drinking water. These results suggest that the model still lacks a thorough understanding of motion and interactions. Leveraging advanced multi-modal large language models to improve the understanding of human-object interactions could be a promising approach to addressing these challenges.
@@
We firmly oppose the misuse of generative AI for creating harmful content or spreading false information. We do not assume any responsibility for potential misuse by users. Nonetheless, we recognize that our approach, which focuses on animation human images, carries the risk of potential misuse. To address these risks, we are committed to maintaining the highest ethical standards in our research by complying with legal requirements and protecting privacy.